# A Weibull Mixture Cure Frailty Model for High-dimensional Covariates

### v1

@@ 3.4 GMIFS method for penMCFM
The GMIFS method is a variable selection technique that iteratively enhances the predictive power of a model by progressively adding variables, thus it can be adapted to obtain parameter estimations in penMCFM. Unlike traditional forward selection methods, GMIFS enforces a monotonic constraint on the coefficients of the added variables, ensuring that they either remain constant or increase with each iteration. By sequentially introducing variables with controlled increases in their coefficients, GMIFS strikes a balance between model sparsity and predictive accuracy. This method was first introduced by Hastie et al. [38] for solving penalized least squares regression problems. Since then, Makowski and Archer [63] and Hou and Archer [42] provided extensions to deal with Poisson regression and ordinal responses for the high dimensional setting, Yu et al. [104] used it to introduce a variable selection procedure for nonlinear regression, and finally Fu et al. [35] applied it in the context of the Weibull MCM.
In the simulation studies, with the aim of comparing our EM-based method for penMCFM not only with the penalized Weibull MC model introduced in [35], but also with another variable selection algorithm, the same framework established in [35] is employed to adapt the GMIFS method to target the observed likelihood function in (3). For comprehensive details on the implementation, readers are encouraged to refer to Hastie et al. [38] and Fu et al. [35], or to check our code details at https://github.com/fatihki/penMCFM.
@@ 4 Simulation Study
For the sake of performance comparisons in the simulation studies, besides our proposed penMCFM EM-based algorithm (detailed in Sections 3.2-3.3), which we refer to as penMCFM(EM), we also implemented a version of the penMCFM algorithm where we incorporate the GMIFS method (detailed in Section 3.4), named penMCFM(GMIFS), and the GMIFS-based penalized Weibull MC model from Fu et al. [35], named MCM(GMIFS). We are also interested in exploring the performance of a survival model that does not account for a cure fraction. To this aim, the Cox PH model with lasso penalty is employed by using the glmnet [34] R package, where the tuning parameter $\lambda$ that gives the most regularized model such that the cross-validated error is within one standard error of the minimum is chosen. This $\lambda$ value is referred to as “lambda.1se” in [34], and thus we refer to this method as penCox.1se. Note that this method only considers the latency part of our model, namely the one related to the $\bm{\beta}_{p}$ regression coefficients.
@@ 4.1 Data generation under the Weibull MCFM
$F_{pop}^{-1}(u|\mathbf{x,z})=\left\{\theta\alpha^{-1}e^{\mathbf{-\bm{\beta}\mathbf{x}}}\left(\left[1-u/\pi(\mathbf{z})\right]^{-1/\theta}-1\right)\right\}^{1/\gamma}$
@@
       generate $u_{i}\sim U\left(0,1\right)$ compute $\pi(\mathbf{z}_{i})=e^{\mathbf{bz}_{i}}/(1+e^{\mathbf{bz}_{i}})$ generate random survival time $t_{i}^{\ast}$: if $u_{i}<\pi(\mathbf{z}_{i})$ then
@@
      generate censoring time: $c_{i}\sim Exp(\lambda_{c})$ determine survival time and censoring indicator: $t_{i}=\min(c_{i},t_{i}^{\ast})$ and $\delta_{i}=I(t_{i}^{\ast}\leq c_{i})$
@@ 4.3 Performance evaluation
$\widehat{C}=\frac{\sum_{i=1}^{n}\sum_{j=1,i\neq j}^{n}I[\widehat{\bm{\beta}}\mathbf{x}_{i}>\widehat{\bm{\beta}}\mathbf{x}_{j}]I_{i,j}}{\sum_{i=1}^{n}\sum_{j=1,i\neq j}^{n}I_{i,j}},$
where $I_{i,j}=I\left[t_{i}<t_{j},\text{ }\delta_{i}=1\right]+I\left[t_{i}=t_{j},\text{ }\delta_{i}=1,\delta_{j}=0\right]$, and $\widehat{\bm{\beta}}$ is the estimated vector of coefficients. Assano and Hirakawa [3] noted that the value of $\widehat{\bm{\beta}}\mathbf{x}_{i}$ is the same for both cured patients and censored uncured patients, and therefore proposed a modified C-statistics that considers this issue by introducing a cure status weighting for each patient. The weight is defined as $1$ for the uncured patients who experienced the event, 00 for the cured patients, and is equal to the estimate of the uncured probability $\widehat{\pi}(\mathbf{z})$ for the censored patients. Then, the C-statistic is defined by Assano and Hirakawa [3] as
$\widehat{C}_{Cure}=\frac{\sum_{i=1}^{n}\sum_{j=1,i\neq j}^{n}I\left[\widehat{\bm{\beta}}\mathbf{x}_{i}>\widehat{\bm{\beta}}\mathbf{x}_{j}\right]\left\{\vartheta_{j}y_{j}+(1-\vartheta_{j})\widehat{\pi}(\mathbf{z}_{j})\right\}I_{i,j}}{\sum_{i=1}^{n}\sum_{j=1,i\neq j}^{n}\left\{\vartheta_{j}y_{j}+(1-\vartheta_{j})\widehat{\pi}(\mathbf{z}_{j})\right\}I_{i,j}},$ (17)
where $\vartheta_{i}$ is an indicator function taking values $\vartheta_{i}=1$ if $y_{i}=0$ or $1$, and $\vartheta_{i}=0$ if $y_{i}$ is missing, and where $\widehat{\pi}(\mathbf{z}_{j})=e^{\widehat{\mathbf{b}}\mathbf{z}_{j}}/(1+e^{\widehat{\mathbf{b}}\mathbf{z}_{j}})$. Notice that $\widehat{C}_{Cure}$ in (17) reduces to $\widehat{C}$ when $v_{i}=1$ and $y_{i}=1$ for all patients.
@@
Specificity $=$ $\sum_{l=1}^{P}I(\beta_{l}\neq 0\cap\widehat{\beta}_{l}=0)/\sum_{l=1}^{P}I(b_{l}=0)$
@@
Furthermore, according to the model, the probability of being cured is sample-specific: this means that, even when two patients share identical clinical characteristics, their unique genomic profiles can influence the cure probability in different ways. Recall that the uncured probability for the $i$th subject is computed by using the formula $\pi(\mathbf{z}_{i})=e^{\mathbf{bz}_{i}}/(1+e^{\mathbf{bz}_{i}})$, for given $\mathbf{b}$ coefficients and for the $i$th subject covariates $\mathbf{z}_{i}$, $i=1,...,n$. Therefore, the average true uncured probability can be computed as $\sum_{k=1}^{M}\left[\sum_{i=1}^{n}\pi^{k}(\mathbf{z}_{i})/n\right]/M$ over $M$ Monte Carlo runs of the EM algorithm, and over the $n$ subjects in the sample, where $\pi^{k}(\mathbf{z}{}_{i})$ is the true uncured probability for the $i$th subject in the $k$th Monte Carlo simulation, with $M$ being the total number of simulated datasets. Consequently, we can also evaluate the accuracy of model estimates in terms of bias and mean squared error (MSE) of the estimated uncured probability $\widehat{\pi}(\mathbf{z})$, which are calculated as
$Bias\left(\widehat{\pi}(\mathbf{z})\right)=\frac{1}{M}\sum_{k=1}^{M}\left[\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\pi}^{k}(\mathbf{z}_{i})-\pi^{k}(\mathbf{z}_{i})\right\}\right],\text{ }MSE\left(\widehat{\pi}(\mathbf{z})\right)=\frac{1}{M}\sum_{k=1}^{M}\left[\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\pi}^{k}(\mathbf{z}_{i})-\pi^{k}(\mathbf{z}_{i})\right\}^{2}\right],$
@@ 4.4 Simulation Results
From Table 1, it can be observed that our proposed method penMCFM(EM) outperforms other methods in terms of the RME and ERR metrics, except for the most difficult case when $v=0.5$. In almost all scenarios, penMCFM(EM) when choosing $\alpha_{Enet}=0.1$ shows the best prediction and estimation performance for the regression coefficients in the cure part, and therefore we observe similar performance for the estimation of $\pi(\mathbf{z})$.
@@
5 Application to RNA-Seq data from TCGA-BRCA
  • the $147$ genes from the KEGG breast cancer pathway [47],
  • the $70$ genes known as the “MammaPrint” BRCA signatures [18],
When we take the intersection of these sources with our already reduced set of $19938$ RNA-seq features, and combine them with the top $2000$ features, we obtain a total of $2650$ RNA-seq features. Lastly, we filter the samples based on the clinical variables listed in Table 2, resulting in a dataset of $828$ observations and around $89\%$ censoring rate. The remaining portion of the RNA-seq data, which includes $189$ observations and has around $70\%$ censoring rate, will be used as validation data in the validation of the identified genes.
@@ 5.1 Identification of biomarkers
To identify the relevant features (biomarkers), we initially partition the dataset randomly into a $80\%$ training and a $20\%$ testing set. In order to select the tuning parameter $\lambda_{Enet}$, we employ a $4$-fold cross-validation on the training dataset to obtain the optimal $\lambda_{Enet}$ that maximizes the C-statistic, $\widehat{C}_{Cure}$ or $\widehat{C}$. We additionally tune the $\alpha_{Enet}$ parameter, by considering values in the set $\{0.1,0.3,0.5,0.7,0.9,1\}$, as the one maximizing the C-statistic value in the testing dataset when employing the $\lambda_{Enet}$ parameter selected from the training dataset. We then repeat this random splitting (and subsequent parameter tuning) $20$ times to get several new training and testing datasets, as a part of our experiments. It is crucial to emphasize the importance of such approach to ensure feature selection robustness, as the selected features are consolidated as the union of the genes associated with non-zero coefficients in all $20$ experiments.
@@
We employ our methods in both Scenarios 1 and 2 as detailed above. The $\alpha_{Enet}$ tuning parameter is determined based on the average values of the C-statistic over the $20$ repeats for the penMCFM(EM) and penCox.1se methods, the only ones allowing this parameter. The results obtained using the optimally selected $\alpha_{Enet}$ parameters can be seen in Figure 5, which shows the boxplots of the C-statistic for the different methods. The number of identified biomarker genes is different for the different methods: for instance, in Scenario $2$, the penMCFM(EM) method with $\alpha_{Enet}=1$ selects $182$ genes for the latency part of the model, genes which correspondingly show nonzero values in $\bm{\beta}_{p}$, while MCM(GMIFS) selects $261$ genes. These identified sets of biomarker genes show great overlap across different methods, as shown in Figure 6, which illustrates the overlap of the selected genes for the latency part of the model across the four methods. Interestingly, penMCFM(EM) generally does not select any features for the incidence part of the model, meaning that no coefficient in $\textbf{b}_{p}$ is estimated to be significantly different from zero, while the GMIFS method (when used with both models penMCFM and MCM) usually selects some features. A similar result was also observed in Fu et al. [35] in an application to acute myeloid leukemia data. A summary of the overlap of the selected genes of these two methods for the incidence part of the model is also given in Figure S1111 of the Supplementary Material. Furthermore, the selected sets of biomarker genes for the different methods, along with their frequencies of occurrence obtained over the $20$ repeats of each method, are provided at https://github.com/fatihki/penMCFM as excel files.
@@
From inspection of Figure 5 one can see that the proposed method penMCFM(EM) achieves the second-highest C-statistic values, whereas MCM(GMIFS) attains the highest C-statistic. It is worth noting however that the latter method selects a larger number of genes for both parts of the model compared to other methods, which may contribute to its higher C-statistic values. For a fair comparison of the methods performance, we carry out more investigations for the identified biomarker genes via gene set enrichment analyses, and by constructing a prognostic risk score, as described in Sections 5.2 and 5.3, respectively. Furthermore, we validate a subset of the top-identified genes derived from penMCFM(EM) when considering the intersection of results from both scenarios, by referring to the existing BRCA literature. This additional validation can be found in Table S3333 of the Supplementary Material.
@@
Figure 5: Results of the analysis of the TCGA-BRCA data. Boxplots of the C-statistic values obtained on the testing datasets over the $20$ repeated data-splitting processes

### v2

@@ 3.4 GMIFS method for penMCFM
The GMIFS method is a variable selection technique that iteratively enhances the predictive power of a model by progressively adding variables, thus it can be adapted to obtain parameter estimations in penMCFM. Unlike traditional forward selection methods, GMIFS enforces a monotonic constraint on the coefficients of the added variables, ensuring that they either remain constant or increase with each iteration. By sequentially introducing variables with controlled increases in their coefficients, GMIFS strikes a balance between model sparsity and predictive accuracy. This method was first introduced by Hastie et al. [36] for solving penalized least squares regression problems. Since then, Makowski and Archer [39] and Hou and Archer [40] provided extensions to Poisson regression and ordinal responses in the high dimensional setting, Yu et al. [41] to variable selection for nonlinear regression, and Fu et al. [35] to the Weibull MCM.
In the simulation studies, with the aim of comparing our EM-based method for penMCFM not only with the penalized Weibull MC model introduced in [35], but also with another variable selection algorithm, the GMIFS method is adapted so to be able to target our observed likelihood function in (6). For a comprehensive introduction to this method, readers are encouraged to refer to Hastie et al. [36] and Fu et al. [35], and for further details on the implementation, to check our code details at https://github.com/fatihki/penMCFM.
@@ 4 Simulation Study
For the sake of performance comparisons in the simulation studies, besides our proposed penMCFM EM-based algorithm (detailed in Sections 3.2-3.3), which we refer to as penMCFM(EM), we also implemented a version of the penMCFM algorithm where we incorporate the GMIFS method (detailed in Section 3.4), named penMCFM(GMIFS), and the GMIFS-based penalized Weibull MC model from Fu et al. [35], named MCM(GMIFS). We are also interested in exploring the performance of a survival model that does not account for a cure fraction. To this aim, the Cox PH model with lasso penalty is employed by using the glmnet [16] R package based on the observed data $(\bm{\delta},\mathbf{t},\mathbf{x_{u}},\mathbf{x_{p}})$, where the tuning parameter $\lambda$ that gives the most regularized model such that the cross-validated error is within one standard error of the minimum is chosen. This $\lambda$ value is referred to as “lambda.1se” in [16], and thus we refer to this method as penCox.1se. Note that this method only considers the latency part of our model, namely the one related to the $\bm{\beta}_{p}$ regression coefficients.
@@ 4.1 Data generation under the Weibull MCFM
$F_{pop}^{-1}(u|\mathbf{x,z})=\left\{\theta\alpha^{-1}e^{\mathbf{-\mathbf{x}^{\top}\bm{\beta}}}\left(\left[1-u/\pi(\mathbf{z})\right]^{-1/\theta}-1\right)\right\}^{1/\gamma}$
@@
      generate $u_{i}\sim U\left(0,1\right)$
      compute $\pi(\mathbf{z}_{i})=e^{\mathbf{z}_{i}^{\top}\mathbf{b}}/(1+e^{\mathbf{z}_{i}^{\top}\mathbf{b}})$
      generate random survival time $t_{i}^{\ast}$:
      if $u_{i}<\pi(\mathbf{z}_{i})$ then
@@
      generate censoring time: $c_{i}\sim Exp(\lambda_{c})$
      determine survival time and censoring indicator:
       $t_{i}=\min(c_{i},t_{i}^{\ast})$ and $\delta_{i}=I(t_{i}^{\ast}\leq c_{i})$
@@ 4.3 Performance evaluation
$\widehat{C}=\frac{\sum_{i=1}^{n}\sum_{j=1,i\neq j}^{n}I(\mathbf{x}_{i}^{\top}\widehat{\bm{\beta}}>\mathbf{x}_{j}^{\top}\widehat{\bm{\beta}})I_{i,j}}{\sum_{i=1}^{n}\sum_{j=1,i\neq j}^{n}I_{i,j}},$
where $I_{i,j}=I[t_{i}<t_{j},\delta_{i}=1]+I[t_{i}=t_{j},\delta_{i}=1,\delta_{j}=0]$, and $\widehat{\bm{\beta}}$ is the estimated vector of coefficients. Assano and Hirakawa [43] noted that the value of $\widehat{\bm{\beta}}\mathbf{x}_{i}$ is the same for both cured patients and censored uncured patients, and therefore proposed a modified C-statistics that considers this issue by introducing a cure status weighting for each patient. The weight is defined as $1$ for the uncured patients who experienced the event, 00 for the cured patients, and is equal to the estimate of the uncured probability $\widehat{\pi}(\mathbf{z})$ for the censored patients. Then, the C-statistic is defined by Assano and Hirakawa [43] as
$\widehat{C}_{Cure}=\frac{\sum_{i=1}^{n}\sum_{j=1,i\neq j}^{n}I(\mathbf{x}_{i}^{\top}\widehat{\bm{\beta}}>\mathbf{x}_{j}^{\top}\widehat{\bm{\beta}})\left\{\vartheta_{j}y_{j}+(1-\vartheta_{j})\widehat{\pi}(\mathbf{z}_{j})\right\}I_{i,j}}{\sum_{i=1}^{n}\sum_{j=1,i\neq j}^{n}\left\{\vartheta_{j}y_{j}+(1-\vartheta_{j})\widehat{\pi}(\mathbf{z}_{j})\right\}I_{i,j}},$ (17)
where $\vartheta_{i}$ is an indicator function taking values $\vartheta_{i}=1$ if $y_{i}=0$ or $1$, and $\vartheta_{i}=0$ if $y_{i}$ is missing, and where $\widehat{\pi}(\mathbf{z}_{j})=e^{\mathbf{z}_{j}^{\top}\widehat{\mathbf{b}}}/(1+e^{\mathbf{z}_{j}^{\top}\widehat{\mathbf{b}}})$. Notice that $\widehat{C}_{Cure}$ in (17) reduces to $\widehat{C}$ when $v_{i}=1$ and $y_{i}=1$ for all patients.
@@
Specificity $=$ $\sum_{l=1}^{P}I(\beta_{l}\neq 0\cap\widehat{\beta}_{l}=0)/\sum_{l=1}^{P}I(\beta_{l}=0)$
@@
Furthermore, according to the model, the probability of being cured is sample-specific: this means that, even when two patients share identical clinical characteristics, their unique genomic profiles can influence the cure probability in different ways. Recall that the uncured probability for the $i$th subject is computed by using the formula $\pi(\mathbf{z}_{i})=e^{\mathbf{z}_{i}^{\top}\mathbf{b}}/(1+e^{\mathbf{z}_{i}^{\top}\mathbf{b}})$, for given $\mathbf{b}$ coefficients and for the $i$th subject covariates $\mathbf{z}_{i}$, $i=1,...,n$. Therefore, the average true uncured probability can be computed as $\sum_{k=1}^{M}\left[\sum_{i=1}^{n}\pi^{k}(\mathbf{z}_{i})/n\right]/M$ over $M$ Monte Carlo runs of the EM algorithm, and over the $n$ subjects in the sample, where $\pi^{k}(\mathbf{z}_{i})$ is the true uncured probability for the $i$th subject in the $k$th Monte Carlo simulation, with $M$ being the total number of simulated datasets. Consequently, we can also evaluate the accuracy of model estimates in terms of bias and mean squared error (MSE) of the estimated uncured probability $\widehat{\pi}(\mathbf{z})$, which are calculated as
$Bias\left(\widehat{\pi}(\mathbf{z})\right)=\frac{1}{M}\sum_{k=1}^{M}\left[\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\pi}^{k}(\mathbf{z}_{i})-\pi^{k}(\mathbf{z}_{i})\right\}\right],$
and
$MSE\left(\widehat{\pi}(\mathbf{z})\right)=\frac{1}{M}\sum_{k=1}^{M}\left[\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\pi}^{k}(\mathbf{z}_{i})-\pi^{k}(\mathbf{z}_{i})\right\}^{2}\right],$
@@ 4.4 Simulation Results
To assess the accuracy of variable selection, it is vital to consider sensitivity, specificity, and the FPR plots in combination. Figures 1 and 2 reveal that the proposed penMCFM(EM) method exhibits very similar average results when choosing $\alpha_{Enet}$ equal to either values 0.5, 0.9, 10.50.910.5,\>0.9,\>10.5 , 0.9 , 1 across all metrics, and moreover, these figures also show that penMCFM(EM) is the method providing best overall results across all simulated scenarios and considered metrics. Notably, when looking specifically at Figure 1, the penMCFM(GMIFS) method consistently selects the fewest variables, often below the true value of $20$, while conversely the MCM model version, MCM(GMIFS), chooses the largest number of variables compared to other methods. These behaviors obviously reflect in the often suboptimal sensitivity, specificity and FPR values attained by these two methods. The penCox.1se, which utilizes the penalized Cox regression with a lasso penalty, ranks as the second method with regards to the number of selected variables in Figure 1, thus showing large sensitivity but suboptimal specificity and FPR. If we compare the variable selection performance of our proposed penMCFM(EM) method when choosing $\alpha_{Enet}=1$ and penCox.1se, they show similar sensitivity plots although penMCFM(EM) selects less variables than penCox.1se. Moreover, penMCFM(EM) attains better performance across $\alpha_{Enet}$ values than penCox.1se in terms of specificity and FPR across all $\mathbf{v}$ and $\rho$ settings. The MCM (GMIFS) method demonstrates better sensitivity performance for $\mathbf{b}_{p}$ (Figure 2) as compared to $\bm{\beta}_{p}$ (Figure 1). However, similarly to what shown for $\bm{\beta}_{p}$ in Figure 1, penMCFM(EM) consistently outperforms other methods across all $\alpha_{Enet}$ values also for estimating $\mathbf{b}_{p}$ (Figure 2), as it selects the fewest variables and frequently approaches the true value of $20$ more closely than other methods, while maintaining large sensitivity and specificity, and exhibiting low FPR.
@@
From Table 1, it can be observed that our proposed method penMCFM(EM) outperforms other methods in terms of the RME and ERR metrics, except for the most difficult case when $v=0.5$. In almost all scenarios, penMCFM(EM) with $\alpha_{Enet}=0.1$ shows the best prediction and estimation performance for the regression coefficients in the cure part, and therefore we observe similar performance for the estimation of $\pi(\mathbf{z})$.
5 Application to RNA-Seq data from TCGA-BRCA
  • the $147$ genes from the KEGG breast cancer pathway [51],
  • the $70$ genes known as the “MammaPrint” BRCA signatures [53],
When we take the intersection of these sources with our already reduced set of $19938$ RNA-seq features, and combine them with the top $2000$ features, we obtain a total of $2650$ RNA-seq features. Lastly, we filter the samples based on the clinical variables listed in Table 2, resulting in a dataset of $828$ observations and around $89\%$ censoring rate. The remaining portion of the RNA-seq data, which includes $189$ observations and has around $70\%$ censoring rate, will be used for validating the identified genes. Even if the censoring rate is slightly lower on this validation set (as compared to the training and testing sets), filtering on clinical variables makes this validation dataset at most homogeneous to the training and testing sets with respect to covariates, which is most crucial in calculating the Prognostic Risk Score.
@@ 5.1 Identification of biomarkers
To identify the relevant features (biomarkers), we initially partition the dataset randomly into a $80\%$ training and a $20\%$ testing set. In order to select the tuning parameter $\lambda_{Enet}$, we employ a $4$-fold cross-validation on the training dataset to obtain the optimal $\lambda_{Enet}$ that maximizes the C-statistic ($\widehat{C}$ for the penCox.1se method and $\widehat{C}_{Cure}$ for all other methods). We additionally tune the $\alpha_{Enet}$ parameter, by considering values in the set $\{0.1,0.3,0.5,0.7,0.9,1\}$, as the one maximizing the C-statistic value in the testing dataset when employing the $\lambda_{Enet}$ parameter selected from the training dataset. We then repeat this random splitting (and subsequent parameter tuning) $20$ times to get several new training and testing datasets, as a part of our experiments. It is crucial to emphasize the importance of such approach to ensure feature selection robustness, as the selected features are consolidated as the union of the genes associated with non-zero coefficients in all $20$ experiments.
@@
We employ our methods in both Scenarios 1 and 2 as detailed above. The $\alpha_{Enet}$ tuning parameter is determined based on the average values of the C-statistic over the $20$ repeats for the penMCFM(EM) and penCox.1se methods, the only ones allowing this parameter. The results obtained using the optimally selected $\alpha_{Enet}$ parameters can be seen in Figure 5, which shows the boxplots of the C-statistic for the different methods. The number of identified biomarker genes is different for the different methods: for instance, in Scenario $2$, the penMCFM(EM) method with $\alpha_{Enet}=1$ selects $182$ genes for the latency part of the model, genes which correspondingly show nonzero values in $\bm{\beta}_{p}$, while MCM(GMIFS) selects $208$ genes. These identified sets of biomarker genes show great overlap across different methods, as shown in Figure 6, which illustrates the overlap of the selected genes for the latency part of the model across the four methods. Interestingly, penMCFM(EM) generally does not select any features for the incidence part of the model, meaning that no coefficient in $\textbf{b}_{p}$ is estimated to be significantly different from zero, while the GMIFS method (when used with both models penMCFM and MCM) usually selects some features. A similar result was also observed in Fu et al. [35] in an application to acute myeloid leukemia data. A summary of the overlap of the selected genes of these two methods for the incidence part of the model is also given in Figure S1111 of the Supplementary Material. Furthermore, the selected sets of biomarker genes for the different methods, along with their frequencies of occurrence obtained over the $20$ repeats of each method, are provided at https://github.com/fatihki/penMCFM as excel files.
From inspection of Figure 5 one can see that the proposed method penMCFM(EM) achieves the second-largest C-statistic values, whereas MCM(GMIFS) attains the largest C-statistic. It is worth noting however that the latter method selects a larger number of genes for both parts of the model compared to other methods, which may contribute to its greater C-statistic values. Moreover, it is also possible to consider other metrics, such as the area under the receiver operating characteristic (ROC) curve and the integrated Brier score, for comparing the different methods, for both the incidence and latency parts of the model. To calculate the area under the ROC curve (AUC), each method is used to estimate $5$-years survival for the subjects in the testing set, and compared to outcome on the corresponding time interval. The integrated Brier score (IBS) is a composite measure of discrimination and calibration, and lower values indicate better model calibration and predictive accuracy. The procedures outlined by Zhao et al. [47] were followed to compute both the AUC and IBS metrics. Our analysis leads to the conclusion that the penMCFM(EM) method, as depicted in Figures 5(a) and 5(b), exhibits superior AUC and IBS performance compared to the alternative methods for both the incidence and latency components of the model.
@@
For a fair comparison of the methods performance, we carry out more investigations for the identified biomarker genes via gene set enrichment analyses, and by constructing a prognostic risk score, as described in Sections 5.2 and 5.3, respectively. Furthermore, we validate a subset of the top-identified genes derived from penMCFM(EM) when considering the intersection of results from both scenarios, by referring to the existing BRCA literature. This additional validation can be found in Table S3333 of the Supplementary Material.
@@
Figure 5: Results of the analysis of the TCGA-BRCA data. Boxplots of the C-statistic, AUC and IBS values obtained on the testing datasets over the $20$ repeated data-splitting processes